OpenStudy (anonymous):

a certain mountain has elevation of 19,515 feet.in 1915, the glacier on this peak covered 6 acres.by 2002 this glacier had melted to only 1 acre. (a) assume that this glacier melted at a constant rate each year.find this yearly rate. (b) use your awser form part(a) to wite a linear equation that gives the acreage a of this glacier t years past 1915. the yearly rate of change is__acres/year. what is the equation that gives the acreage of the glacier t yaers after 1915?

OpenStudy (anonymous):

im lost

OpenStudy (anonymous):

use the points (0, 6) and (87, 1) to find the slope, or m, also the yearly rate of change. I got 87 because that is how many years elapsed from 1915 to 2002. If the equation has to be in standard form, use y=mx+b. You already will have solved for m, and then use the first x,y coordinate to determine b. This will give you the equation

OpenStudy (anonymous):

(87/2,7/2)

OpenStudy (anonymous):

are you thinking of the midpoint formula? Slope would be (y2-y1)/(x2-x1)

OpenStudy (anonymous):

im lost

OpenStudy (anonymous):

ok it 87 acres

OpenStudy (anonymous):

a year

OpenStudy (anonymous):

slope=m= (1-6)/ (87-0) = -5/87 acres/year which intuitively makes sense that its negative because the glacier size is decreasing. This was derived by finding the change in y (5 acres over that period of time) divided by the elapsed time (87 years)

OpenStudy (anonymous):

5/87 acres/year

OpenStudy (anonymous):

actually \(-\frac{5}{87}\) since they are decreasing

OpenStudy (anonymous):

so if \(x\) is the number of years after 1915 then the equation would be \[y=-\frac{5}{87}x+6\]

OpenStudy (anonymous):

the reason for the \(6\) out at the end is because that is what you started with i.e. if \(x=0\) then \(y=6\)